OpenStudy (anonymous):

Need help fast! Variable y varies directly with x^2, and y = 96 when x = 4. What is the value of y when x = 2? y = ?

OpenStudy (anonymous):

@Hero @Luigi0210

OpenStudy (luigi0210):

For this specific problem your equation would be \(y=kx^2 \) Plug in the given values to solve for k

OpenStudy (anonymous):

Um how do i find k? isnt k=y/x? So... 384? that couldnt work...

OpenStudy (anonymous):

A little help? @Luigi0210

OpenStudy (anonymous):

@Hero @bohotness @hhelpplzzzz @k_lynn

OpenStudy (mathstudent55):

No. \(y = kx^2\) is the general form of quadratic variation. A specific case can be \(y = 2x^2\) or \(y = \dfrac{2}{3}x^2\). What you need to do is to find what is the value of k in your specific problem.

OpenStudy (mathstudent55):

Start with the general form: \(y = kx^2\) We need to find the value of k, but there are 3 unknowns in the equation. The problem tells you that when x = 4, y = 96. You use the general equation and plug in 4 for x and 96 for y. Now the only unknown is k. You can now solve for k. Once you find y, you can rewrite the equation with the correct value of k plugged in for k.

OpenStudy (anonymous):

Yeah but... how do i find k? and in this case dosent it mean the awnser is 96? @mathstudent55

OpenStudy (anonymous):

96 diveded by 2*

OpenStudy (mathstudent55):

No. Here it is step by step. 

OpenStudy (mathstudent55):

Since the problem tells us y = 96 when x = 4, we replace x with 4 and y with 96.